OpenStudy (anonymous):

someone please help me solve these using quadratic formula and please explain :) a. 1=8x (x+1) b. v^2+4v-60 c. x^2+2radical3x+3=0

OpenStudy (anonymous):

sorryyyyyyyyyyyyyyyyyyyyyyyyyyyyyyy

OpenStudy (mimi_x3):

a) 1= 8x^2+ 8x 8x^2 + 8x - 1 = 0 a = 8 , b = 8 , c =-1 Use the quadratic formula. b) v^2 + 4v - 60 = 0 (x-6)(x+10) x = 6, -10

OpenStudy (lalaly):